can anyone tell me what are the forms of positive integers? it must be like 2q,2q+1,2q+2.......... now in place of 2 can we take numbers like 4,6,5,22,33,45 etc? also i want to know the forms of prime positive integers .

Grade:10

2 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Sudent,
Please find answer to your question below,


Form of positive integers:
a, a + 1, a + 2, a + 3, …....................
where

a\geq 1

Form of prime number:
n = 6q + 1; 6q + 5
where q is a non negative integer.
It will cover the most ot the prime number.


Sumit Majumdar IIT Delhi
9 years ago
Hello student,
Any positive number should be written as \left | x \right |.
This is the modulus function. The domain can be both positive or negative integer but the range will always be positive integer.
For the given numbers, 2q, 2q+1, … can be negative as well and similarly if we replace 2 by any number it may be positive or negative. So, if a number has to be strictly positive, it should be written as \left | x \right |.
Regards
Sumit